# centertableaux

### v2

@@ Abstract.
Let $G$ be a general linear group over ${\mathbb{R}}$, ${\mathbb{C}}$, or ${\mathbb{H}}$, or a real unitary group. In this paper, we precisely describe the number of isomorphism classes of irreducible Casselman-Wallach representations of $G$ with a given infinitesimal character and a given associated variety, expressed in terms of certain combinatorial data called painted Young diagrams and assigned Young diagrams.
@@ 1.1. Background and goals
In their paper [BMSZ], Barbasch-Ma-Sun-Zhu established a formula for counting the number of isomorphism classes of irreducible representations of a real reductive group $G$ with a given infinitesimal character and associated variety. Using this formula, when $G$ is a classical Lie group, they precisely determined the number of special unipotent representations of $G$ attached to $\check{{\mathcal{O}}}$ in the sense of Arthur and Barbasch-Vogan, where $\check{{\mathcal{O}}}$ is a nilpotent orbit in the Langlands dual group of $G$ (or the metaplectic dual of $G$ when $G$ is the real metaplectic group). Furthermore, they utilized these counting results to construct all special unipotent representations and demonstrated that all such representations are unitarizable.
In this paper, we apply their counting formula to provide a precise description of the number of isomorphism classes of irreducible representations with a given infinitesimal character and associated variety, in the case when $G$ is a general linear group or a unitary group.
@@ 1.2. Representation-theoretic and combinatorial notations
For each Casselman-Wallach representation $V\in{\mathrm{Rep}}(G)$, we define in Section 2.1 an ${\mathrm{Inn}}(\mathfrak{g})$-stable Zariski closed subvariety $\mathrm{AV}(V)$ of ${\mathrm{Nil}}(\mathfrak{g}^{*})$, which is called the complex associated variety of $V$. Moreover, if $V$ is irreducible, we explain that $\mathrm{AV}(V)$ is the closure of a unique nilpotent orbit.
@@ 1.3. The main results
    Suppose $G={\mathrm{GL}}_{3}({\mathbb{R}})$, $\nu=\left(0,1,\frac{1}{2}\right)$. There are two nilpotent orbits ${\mathcal{O}}_{1}$ and ${\mathcal{O}}_{2}$, whose closures can occur as the associated variety of an irreducible Casselman-Wallach representation with this infinitesimal character. The corresponding Young diagrams are, respectively
    {ytableau}&$\ ~{}\quad and\quad\ytableau~{}&~{}~{}.$
    They each have a unique decomposition into small Young diagrams of sizes $2$ and $1$ respectively,
    {ytableau}&={ytableau}$\ ~{}\quad\mathop{\sqcup}\limits^{r}\quad\ytableau~{},$
    {ytableau}&={ytableau}&$\ytableau~{}&~{}~{}\quad=\quad\ytableau~{}&~{}\quad\mathop{\sqcup}\limits^{r}\quad\ytableau~{}.$
    After enumerating all the assignments and paintings, we find that $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}_{1}))=4$ and $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}_{2}))=3$.
@@
    Suppose $G={\mathrm{GL}}_{4}({\mathbb{H}})$, $\nu=(1,1,1,2,2,2,3,3)$. Then the associated variety of irreducible Casselman-Wallach representations with this infinitesimal character can only be the closure of the nilpotent orbit ${\mathcal{O}}$, whose Young diagram is
@@
    Suppose $G={\mathrm{GL}}_{5}({\mathbb{H}})$, $\nu=(1,1,1,1,2,2,2,3,3,4)$. There are two nilpotent orbits ${\mathcal{O}}_{1}$ and ${\mathcal{O}}_{2}$, whose closures can occur as the associated variety of an irreducible Casselman-Wallach representation with this infinitesimal character. The corresponding Young diagrams are, respectively
@@
    (1.10) $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}))=\left\{\begin{aligned} &0,&\textrm{if $\lambda_{1}-\lambda_{1}^{\prime}\notin{\mathbb{Z}}$};\\ &\sharp\left({\mathrm{A}}_{[d_{1},\cdots,d_{k}]}(\iota)\right)\cdot\sharp\left({\mathrm{A}}_{[d_{1}^{\prime},\cdots,d_{l}^{\prime}]}(\iota^{\prime})\right)\cdot\delta_{\iota,\iota^{\prime}},&\textrm{if $\lambda_{1}-\lambda_{1}^{\prime}\in{\mathbb{Z}}$}\end{aligned}\right.$
@@
    $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}))=$ $\left(\sum_{\begin{subarray}{c}(\iota,\iota^{\prime},\iota_{1},\iota_{1}^{\prime}\cdots,\iota_{r},\iota_{r}^{\prime})\in\mathrm{YD}_{e}\times\mathrm{YD}_{e^{\prime}}\times\mathrm{YD}_{e_{1}}\times\mathrm{YD}_{e_{1}}\times\cdots\times\mathrm{YD}_{e_{r}}\times\mathrm{YD}_{e_{r}}\\ \iota\mathop{\sqcup}\limits^{r}\iota^{\prime}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\cdots\mathop{\sqcup}\limits^{r}\iota_{r}\mathop{\sqcup}\limits^{r}\iota_{r}=\iota({\mathcal{O}})\end{subarray}}\sharp({\mathrm{Irr}}_{{\boldsymbol{\lambda}}}({\mathrm{U}}(p^{\prime},q^{\prime});{\mathcal{O}}(\iota^{\prime})))\right)$
@@
    $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}))=$ $\left(\sum_{\begin{subarray}{c}(\iota,\iota^{\prime},\iota_{1},\iota_{1}^{\prime}\cdots,\iota_{r},\iota_{r}^{\prime})\in\mathrm{YD}_{e}\times\mathrm{YD}_{e^{\prime}}\times\mathrm{YD}_{e_{1}}\times\mathrm{YD}_{e_{1}}\times\cdots\times\mathrm{YD}_{e_{r}}\times\mathrm{YD}_{e_{r}}\\ \iota\mathop{\sqcup}\limits^{r}\iota^{\prime}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\cdots\mathop{\sqcup}\limits^{r}\iota_{r}\mathop{\sqcup}\limits^{r}\iota_{r}=\iota({\mathcal{O}})\end{subarray}}\sharp({\mathrm{Irr}}_{{\boldsymbol{\lambda}}}({\mathrm{U}}(p^{\prime},q^{\prime});{\mathcal{O}}(\iota^{\prime})))\right)$
@@ 2. Preliminaries
In this section, we review some preliminary material relevant to the counting formula, following the exposition in [BMSZ].
2.1. Associated varieties and genuine representations
@@
is automatically a ${\mathcal{Z}}(\mathfrak{g})$-finite two-sided ideal, we define the complex associated variety $\mathrm{AV}(V)$ of $V$ as the associated variety of the ideal ${\mathrm{Ann}}(V)$, which is a ${\mathrm{Inn}}(\mathfrak{g})$-stable closed subvariety of ${\mathrm{Nil}}(\mathfrak{g}^{*})$. Suppose $S$ is a ${\mathrm{Inn}}(\mathfrak{g})$-stable Zariski closed subset of ${\mathrm{Nil}}(\mathfrak{g}^{*})$, denote by ${\mathrm{Rep}}_{S}(G)$ the category of Casselman-Wallach representations of $G$ whose complex associated variety is contained in $S$.
By the work of Borho and Brylinski [BB], and Joseph [Jos], the associated variety of any primitive ideal (annihilator of an irreducible $\mathfrak{g}$-module) is the Zariski closure of a single nilpotent orbit. This implies the associated variety of any irreducible Casselman-Wallach representation is the Zariski closure of a single nilpotent orbit.
Now we assume $G=\widetilde{{\mathrm{U}}}(p,q)$ for the remainder of this subsection. A representation of $\widetilde{{\mathrm{U}}}(p,q)$ is called genuine if the central subgroup {±$\{\pm 1\}$, which is the kernel of the covering homomorphism $\widetilde{{\mathrm{U}}}(p,q)\to{\mathrm{U}}(p,q)$ acts on it through the nontrivial character. Let ${\mathrm{Rep}}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q)\right)$ and ${\mathrm{Rep}}_{\nu}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q)\right)$ denote the full subcategory of genuine representations and the full subcategory of genuine representations with infinitesimal character $\nu$, respectively. Their Grothendieck groups are denoted by ${\mathcal{K}}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q)\right)$ and ${\mathcal{K}}_{\nu}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q)\right)$, respectively. The irreducible objects in these categories are denoted by ${\mathrm{Irr}}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q)\right)$ and ${\mathrm{Irr}}_{\nu}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q)\right)$, respectively.
Let ${\mathrm{Irr}}_{\nu}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q);{\mathcal{O}}\right)$ denote the subset of ${\mathrm{Irr}}_{\nu}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q)\right)$ consisting of representations with associated variety $\overline{{\mathcal{O}}}$.
Since we have a bijection
@@
the number of genuine irreducible representations of $\widetilde{{\mathrm{U}}}(p,q)$ with infinitesimal character $\nu$ and associated variety $\overline{{\mathcal{O}}}$ is the same as the number of irreducible representations of ${\mathrm{U}}(p,q)$ with infinitesimal character $\nu+\mu$ and associated variety $\overline{{\mathcal{O}}}$, where $\mu\in{{}^{a}\mathfrak{h}^{*}}$ is the only element in the $W$-orbit corresponds to the infinitesimal character of $\mathrm{det}^{\frac{1}{2}}$.
@@ 3.4.2. General case
$[d_{i,1},\cdots,d_{i,k_{i}}]$ is a partition of $e_{i}$, $[d_{j,1}^{\prime},\cdots,d_{j,l_{j}}^{\prime}]$ is a partition of $e_{j}^{\prime}$, and $\lambda_{i,p}-\lambda_{i,q}\in{\mathbb{Z}}$, $\lambda_{i,1}-\lambda_{j,1},\ \lambda_{i,1}^{\prime}-\lambda_{j,1}^{\prime}\notin{\mathbb{Z}}$ for different $i,j$. The integral Weyl group is $W_{\Lambda}=W(\Lambda)=(\mathrm{S}_{e_{1}}\times\cdots\mathrm{S}_{e_{r}})\times(\mathrm{S}_{e_{1}}\times\cdots\times\mathrm{S}_{e_{r}})\subseteq\mathrm{S}_{n}\times\mathrm{S}_{n}=W$.
@@
(3.32) $\varphi:$ ${\mathcal{P}}_{\Lambda_{1}}({\mathrm{G}}_{A^{{\mathbb{C}}}}(V_{1}))\times\cdots\times{\mathcal{P}}_{\Lambda_{r}}({\mathrm{G}}_{A^{{\mathbb{C}}}}(V_{r}))$$\longrightarrow$ $\quad{\mathcal{P}}_{\Lambda}(G)$
@@
(3.35) $=\sum_{\begin{subarray}{c}(\iota_{1},\cdots,\iota_{r}),(\iota_{1}^{\prime},\cdots,\iota_{r}^{\prime})\in\mathrm{YD}_{e_{1}}\times\cdots\times\mathrm{YD}_{e_{r}}\\ \iota_{1}\mathop{\sqcup}\limits^{r}\cdots\mathop{\sqcup}\limits^{r}\iota_{r}=\iota,\ \iota_{1}^{\prime}\mathop{\sqcup}\limits^{r}\cdots\mathop{\sqcup}\limits^{r}\iota_{r}^{\prime}=\iota^{\prime}\end{subarray}}\prod_{i=1}^{r}\sharp\left({\mathrm{Irr}}_{{\boldsymbol{\lambda}}_{i}}({\mathrm{G}}_{A^{{\mathbb{C}}}}(V_{i});{\mathcal{O}}_{\iota_{i}}\times{\mathcal{O}}_{\iota_{i}^{\prime}})\right).$

### v3

@@ Abstract.
Let $G$ be a general linear group over ${\mathbb{R}}$, ${\mathbb{C}}$, or ${\mathbb{H}}$, or a real unitary group. In this paper, we precisely describe the number of isomorphism classes of irreducible Casselman-Wallach representations of $G$ with a given infinitesimal character and a given complex associated variety, expressed in terms of certain combinatorial data called painted Young diagrams and assigned Young diagrams.
@@ 1.1. Background and goals
Let $G$ be a real reductive Lie group. Due to the work of Harish-Chandra, the set ${\mathrm{Irr}}(G)$ of infinitesimally equivalence classes of irreducible admissible representations (equivalently, isomorphism classes of Casselman-Wallach representations [Wal, Chapter 11]) of $G$ admit a partition according to infinitesimal characters:
${\mathrm{Irr}}(G)=\bigsqcup_{\lambda}{\mathrm{Irr}}_{\lambda}(G),$
where each subset ${\mathrm{Irr}}_{\lambda}(G)$ is finite ([Har, Theorem 7]), the number of irreducible representations in ${\mathrm{Irr}}_{\lambda}(G)$ can be effectively computed by listing the corresponding Langlands parameters ([ALTV, Section 6]).
Another invariant of irreducible Casselman-Wallach representations is the complex associated variety, which turns out to be the closure of a unique nilpotent orbit. This notion plays an important role in several areas, including Kirillov’s orbit philosophy and the Kazhdan-Lusztig theory.
In this paper, we study the further partition of each ${\mathrm{Irr}}_{\lambda}(G)$ according to the complex associated varieties:
${\mathrm{Irr}}_{\lambda}(G)=\bigsqcup_{{\mathcal{O}}}{\mathrm{Irr}}_{\lambda}(G;{\mathcal{O}}),$
and we give a precise description of the size of each ${\mathrm{Irr}}_{\lambda}(G;{\mathcal{O}})$ when $G$ is a general linear group over ${\mathbb{R}}$, ${\mathbb{C}}$, or ${\mathbb{H}}$, or a real unitary group.
The counting method employed in this paper was developed in [BMSZ] by Barbasch, Ma, Sun, and Zhu, based on the theory of coherent continuation representations. This theory, which first appeared in [BV83], will be introduced in Section 2.
@@ 1.2. Representation-theoretic and combinatorial notations
For each Casselman-Wallach representation $V\in{\mathrm{Rep}}(G)$, we define in Section 2.1 an ${\mathrm{Inn}}(\mathfrak{g})$-stable Zariski closed subvariety $\mathrm{AV}_{\mathbb{C}}(V)$ of ${\mathrm{Nil}}(\mathfrak{g}^{*})$, which is called the complex associated variety of $V$. Moreover, if $V$ is irreducible, we explain that $\mathrm{AV}_{\mathbb{C}}(V)$ is the closure of a unique nilpotent orbit.
@@ 1.3. The main results
    Suppose $G={\mathrm{GL}}_{n}({\mathbb{R}})$, recall that an irreducible Casselman-Wallach representation is called generic if it admits a Whittaker model or equivalently its complex associated variety is the closure of principle orbit ([Vog78]), whose Young diagram is:
[…]
    When $\nu=(\lambda_{1},\cdots,\lambda_{n})$ is regular and integral, which means $\0\lambda_{i}-\lambda_{j}\in\mathbb{Z}\backslash\{0\}$ for all $i\neq j$, there is only one assignment of type $[1,1,\cdots,1]$, and exactly $n+1$ paintings of type $A_{{\mathbb{R}}}$ on this Young diagram. Consequently, the number of generic representation of regular integral infinitesimal character is $n+1$.
@@
    Suppose $G={\mathrm{GL}}_{4}({\mathbb{H}})$, $\nu=(1,1,1,2,2,2,3,3)$. Then the complex associated variety of irreducible Casselman-Wallach representations with this infinitesimal character can only be the closure of the nilpotent orbit ${\mathcal{O}}$, whose Young diagram is
@@
    Suppose $G={\mathrm{GL}}_{5}({\mathbb{H}})$, $\nu=(1,1,1,1,2,2,2,3,3,4)$. There are two nilpotent orbits ${\mathcal{O}}_{1}$ and ${\mathcal{O}}_{2}$, whose closures can occur as the complex associated variety of an irreducible Casselman-Wallach representation with this infinitesimal character. The corresponding Young diagrams are, respectively
@@
    (1.10) $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}))=\left\{\begin{aligned} &\sharp\left({\mathrm{A}}_{[d_{1},\cdots,d_{k}]}(\iota)\right)\cdot\sharp\left({\mathrm{A}}_{[d_{1}^{\prime},\cdots,d_{l}^{\prime}]}(\iota^{\prime})\right)\cdot\delta_{\iota,\iota^{\prime}},&\textrm{if $\lambda_{1}-\lambda_{1}^{\prime}\in{\mathbb{Z}}$};\\ &0,&\textrm{if $\lambda_{1}-\lambda_{1}^{\prime}\notin{\mathbb{Z}}$}\end{aligned}\right.$
@@
    $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}))=$ $\sum_{\begin{subarray}{c}(\iota,\iota^{\prime},\iota_{1},\iota_{1}^{\prime}\cdots,\iota_{r},\iota_{r}^{\prime})\in\mathrm{YD}_{e}\times\mathrm{YD}_{e^{\prime}}\times\mathrm{YD}_{e_{1}}\times\mathrm{YD}_{e_{1}}\times\cdots\times\mathrm{YD}_{e_{r}}\times\mathrm{YD}_{e_{r}}\\ \iota\mathop{\sqcup}\limits^{r}\iota^{\prime}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\cdots\mathop{\sqcup}\limits^{r}\iota_{r}\mathop{\sqcup}\limits^{r}\iota_{r}=\iota({\mathcal{O}})\end{subarray}}\sharp({\mathrm{Irr}}_{{\boldsymbol{\lambda}}}({\mathrm{U}}(p^{\prime},q^{\prime});{\mathcal{O}}(\iota^{\prime})))$
@@
    $\sharp({\mathrm{Irr}}_{\nu}(G;{\mathcal{O}}))=$ $\sum_{\begin{subarray}{c}(\iota,\iota^{\prime},\iota_{1},\iota_{1}^{\prime}\cdots,\iota_{r},\iota_{r}^{\prime})\in\mathrm{YD}_{e}\times\mathrm{YD}_{e^{\prime}}\times\mathrm{YD}_{e_{1}}\times\mathrm{YD}_{e_{1}}\times\cdots\times\mathrm{YD}_{e_{r}}\times\mathrm{YD}_{e_{r}}\\ \iota\mathop{\sqcup}\limits^{r}\iota^{\prime}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\cdots\mathop{\sqcup}\limits^{r}\iota_{r}\mathop{\sqcup}\limits^{r}\iota_{r}=\iota({\mathcal{O}})\end{subarray}}\sharp({\mathrm{Irr}}_{{\boldsymbol{\lambda}}}({\mathrm{U}}(p^{\prime},q^{\prime});{\mathcal{O}}(\iota^{\prime})))$
@@ 2. Preliminaries
2.1. Associated varieties
@@
is automatically a ${\mathcal{Z}}(\mathfrak{g})$-finite two-sided ideal, we define the complex associated variety $\mathrm{AV}_{\mathbb{C}}(V)$ of $V$ as the associated variety of the ideal ${\mathrm{Ann}}(V)$, which is a ${\mathrm{Inn}}(\mathfrak{g})$-stable closed subvariety of ${\mathrm{Nil}}(\mathfrak{g}^{*})$. Suppose $S$ is a ${\mathrm{Inn}}(\mathfrak{g})$-stable Zariski closed subset of ${\mathrm{Nil}}(\mathfrak{g}^{*})$, denote by ${\mathrm{Rep}}_{S}(G)$ the category of Casselman-Wallach representations of $G$ whose complex associated variety is contained in $S$.
By the work of Borho and Brylinski [BB], and Joseph [Jos], the associated variety of any primitive ideal (annihilator of an irreducible $\mathfrak{g}$-module) is the Zariski closure of a single nilpotent orbit. This implies the complex associated variety of any irreducible Casselman-Wallach representation is the Zariski closure of a single nilpotent orbit.
2.2. Genuine representations of $\widetilde{{\mathrm{U}}}(p,q)$
In this section we assume $G=\widetilde{{\mathrm{U}}}(p,q)$. A representation of $\widetilde{{\mathrm{U}}}(p,q)$ is called genuine if the central subgroup {±$\{\pm 1\}$, which is the kernel of the covering homomorphism $\widetilde{{\mathrm{U}}}(p,q)\to{\mathrm{U}}(p,q)$ acts on it through the nontrivial character. Let ${\mathrm{Rep}}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q)\right)$ and ${\mathrm{Rep}}_{\nu}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q)\right)$ denote the full subcategory of genuine representations and the full subcategory of genuine representations with infinitesimal character $\nu$, respectively. Their Grothendieck groups are denoted by ${\mathcal{K}}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q)\right)$ and ${\mathcal{K}}_{\nu}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q)\right)$, respectively. The irreducible objects in these categories are denoted by ${\mathrm{Irr}}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q)\right)$ and ${\mathrm{Irr}}_{\nu}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q)\right)$, respectively.
Let ${\mathrm{Irr}}_{\nu}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q);{\mathcal{O}}\right)$ denote the subset of ${\mathrm{Irr}}_{\nu}^{\mathrm{gen}}\left(\widetilde{{\mathrm{U}}}(p,q)\right)$ consisting of representations with complex associated variety $\overline{{\mathcal{O}}}$.
Note that there is a bijection
@@ 2.2. Genuine representations of $\widetilde{{\mathrm{U}}}(p,q)$
where $\mu\in{{}^{a}\mathfrak{h}^{*}}$ is the only element in the $W$-orbit corresponds to the infinitesimal character of $\mathrm{det}^{\frac{1}{2}}$. Therefore, the number of genuine irreducible representations of $\widetilde{{\mathrm{U}}}(p,q)$ with infinitesimal character $\nu$ and complex associated variety $\overline{{\mathcal{O}}}$ is the same as the number of irreducible representations of ${\mathrm{U}}(p,q)$ with infinitesimal character $\nu+\mu$ and complex associated variety $\overline{{\mathcal{O}}}$.
@@ 2.3. Generalities on coherent families of Casselman-Wallach representations
This section is devoted to the introduction of coherent continuation representations, as presented in [BMSZ].
@@ 3.4.2. General case
$[d_{i,1},\cdots,d_{i,k_{i}}]$ is a partition of $e_{i}$, $[d_{j,1}^{\prime},\cdots,d_{j,l_{j}}^{\prime}]$ is a partition of $e_{j}^{\prime}$, and $\lambda_{i,p}-\lambda_{i,q}\in{\mathbb{Z}}$, $\lambda_{i,1}-\lambda_{j,1},\ \lambda_{i,1}^{\prime}-\lambda_{j,1}^{\prime}\notin{\mathbb{Z}}$ for different $i,j$. The integral Weyl group is $W_{\Lambda}=W(\Lambda)=(\mathrm{S}_{e_{1}}\times\cdots\times\mathrm{S}_{e_{r}})\times(\mathrm{S}_{e_{1}}\times\cdots\times\mathrm{S}_{e_{r}})\subseteq\mathrm{S}_{n}\times\mathrm{S}_{n}=W$.
@@
(3.32) $\varphi:$ ${\mathcal{P}}_{\Lambda_{1}}({\mathrm{G}}_{A^{{\mathbb{C}}}}(V_{i}))\times\cdots\times{\mathcal{P}}_{\Lambda_{r}}({\mathrm{G}}_{A^{{\mathbb{C}}}}(V_{r}))$$\longrightarrow$ $\quad{\mathcal{P}}_{\Lambda}(G)$
@@
(3.35) $=\sum_{\begin{subarray}{c}(\iota_{1},\cdots,\iota_{r}),(\iota_{1}^{\prime},\cdots,\iota_{r}^{\prime})\in\mathrm{YD}_{e_{1}}\times\cdots\times\mathrm{YD}_{e_{r}}\\ \iota_{1}\mathop{\sqcup}\limits^{r}\cdots\mathop{\sqcup}\limits^{r}\iota_{r}=\iota,\ \iota_{1}^{\prime}\mathop{\sqcup}\limits^{r}\cdots\mathop{\sqcup}\limits^{r}\iota_{r}^{\prime}=\iota^{\prime}\end{subarray}}\prod_{i=1}^{r}\sharp\left({\mathrm{Irr}}_{{\boldsymbol{\lambda}}_{i}}({\mathrm{GL}}_{e_{i}}({\mathbb{C}});{\mathcal{O}}_{\iota_{i}}\times{\mathcal{O}}_{\iota_{i}^{\prime}})\right).$